Moore Corporation bases its predetermined overhead rate on the estimated machine-hours for the upcoming year. Data for the most recently completed year appear below:

Estimates made at the beginning of the year:
Estimated machine-hours 20,200
Estimated variable manufacturing overhead $ 9.00 per machine-hour
Estimated total fixed manufacturing overhead $ 496,920
Actual machine-hours for the year 21,400


The predetermined overhead rate for the recently completed year was closest to:

$33.60.

$9.00.

$24.60.

$32.22.
Answer #1

Total estimated overhead=Total estimated variable overhead+Total fixed overhead estimated

=(20200*9)+496920=$678720

predetermined overhead rate=Total estimated overhead/Estimated machine hours

=678720/20200

=$33.6